# Online Flow Time Minimization: Tight Bounds for Non-Preemptive Algorithms

### v2

@@ 6 Deterministic Algorithm with Kill-and-Restart
In the absence of blocked small jobs and retired large jobs, the flow time of all jobs in $S$ would be bounded by $F(\textup{{NSJF}}(S))$, which is at most $O\left(\sqrt{n/m}\right)\cdot F(\textup{{OPT}}(S))$ by Lemma 2. The reason is that the processing of active large jobs never affects the scheduling of unblocked small jobs. Whenever NSJF wants to schedule an unblocked small job $j$, the algorithm cannot have a large job smaller than $j$ outside of $S$. If such a large job is active, it must be larger than all small jobs by our online partition rule. If it is retired and still waiting, it is proxied and, therefore, never considered for scheduling.
However, the presence of blocked small jobs and retired large jobs may delay the scheduling of unblocked small jobs, because we mix them together. We define a set $U$, consisting of the retired large jobs, and recall the set $B$, consisting of the blocked small jobs. Correspondingly, we construct two modified instances $U^{\prime}$ and $B^{\prime}$ by updating the release times of all jobs $j\in U$ and $j\in B$ to $s_{j}$; note that our algorithm’s schedule on $S\cup U\cup B$ is identical to that of $\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime})$. By Lemma 2, we know that
$F(\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime}))\leq F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))+O(n\tau)\leq F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))+O(\sqrt{n/m})\cdot\textup{{OPT}}.$
This alone does not prove the lemma, because the jobs in $U^{\prime}$ and $B^{\prime}$ may have later release times than in the original instance, so $F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))$ is not necessarily upper bounded by OPT. However, for every retired job $j\in U^{\prime}$, by the same reason as Lemma 1, it is retired after $s_{j}$, and we can find a newly released job $j^{\prime}$, where $p_{j^{\prime}}>p_{j}$, and $r_{j^{\prime}}>s_{j}$, where $s_{j}$ is the release time of $j$ in $U^{\prime}$. So, at least $S\cup U^{\prime}$ is a weaker set of $J$, and $F(\textup{{OPT}}(S\cup U^{\prime}))\leq\textup{{OPT}}$.
We now aim to bound the difference between $F(\textup{{OPT}}(S\cup U^{\prime}\cup B^{\prime}))$ and $F(\textup{{OPT}}(S\cup U^{\prime}))$. To do so, we partition the timeline into intervals free of unproxied large jobs, denoted as $t_{h},t_{h+1})$ for $h=1,2,\dots$, where every $t_{h}$ can be viewed as the time we start a currently unproxied large job (it may be killed or retired later). In each interval $t_{h},t_{h+1})$, no unproxied large job is processed. Unblocked small jobs are only released in such intervals.
Let $S_{h}$ be the set of unblocked small jobs with $r_{j}\in[t_{h},t_{h+1})$ and retired large jobs who becomes small in $t_{h},t_{h+1})$. Clearly, $\bigcup_{h}S_{h}=S\cup U^{\prime}$ by definition. We similarly define $B^{\prime}_{h}\subseteq B^{\prime}$ to be the jobs in $B^{\prime}$ whose processing falls within $t_{h},t_{h+1})$. Observe that, because only $B^{\prime}_{h}$ can affect $S_{h}$, we have
$\sum_{j\in S_{h}}F_{j}(\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime}))=\sum_{j\in S_{h}}F_{j}(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h}))\leq F(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h})).$
An important property of our algorithm is that $|B^{\prime}_{h}|\leq\sqrt{nm}$; otherwise, we would have killed all unproxied jobs before time $t_{h}$. We will use this property to bound $F(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h})$.
@@
Now consider an optimal preemptive solution $\textup{{OPT}}(S_{h})$. We insert the jobs in $B^{\prime}_{h}$ into $\textup{{OPT}}(S_{h})$ one by one, in a migratory fashion: for each job $j\in B^{\prime}_{h}$, we scan from $r_{j}$ and process it at the first available idle slot, without delaying the original schedule. If $j$’s processing overlaps a busy period, we interrupt it and resume at the next idle slot, possibly on a different machine.
Let $F^{\prime}_{j}$ be the flow time of such a job $j$. It consists of $p_{j}$ plus some waiting time. This waiting time falls entirely in busy periods and is thus upper bounded by $\sum_{i\in S_{h}\cup B^{\prime}_{h}}p_{i}/m$. Since there are at most $\sqrt{nm}$ jobs in $B^{\prime}_{h}$ (by our restart rule), we get:
$F(\textup{{OPT}}^{m}(S_{h}\cup B^{\prime}_{h}))\leq F(\textup{{OPT}}(S_{h}))+O(\sqrt{n/m})\sum_{j\in S_{h}\cup B^{\prime}_{h}}p_{j}.$
Here, we use $\textup{{OPT}}^{m}$ to denote the optimal preemptive solution with migration. Note that migration is required in the above construction, and NSJF is compared against $\textup{{OPT}}^{m}$ in Lemma 2. Summing over all $h$, we obtain:
$\sum_{j\in S}F_{j}(\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime}))$ $\leq\sum_{h}F(\textup{{NSJF}}(S_{h}\cup B^{\prime}_{h}))$
$\leq\sum_{h}F(\textup{{OPT}}^{m}(S_{h}\cup B^{\prime}_{h}))+O\left(\frac{|S_{h}|+|B^{\prime}_{h}|}{\sqrt{nm}}\right)\cdot\textup{{OPT}}$
$\leq\sum_{h}F(\textup{{OPT}}(S_{h}))+O\left(\frac{|S_{h}|+|B^{\prime}_{h}|}{\sqrt{nm}}\right)\cdot\textup{{OPT}}+O(\sqrt{n/m}\sum_{j\in S_{h}\cup B^{\prime}_{h}})p_{j}$
$\leq F(\textup{{OPT}}(S))+O(\sqrt{n/m})\cdot\textup{{OPT}}+O(\sqrt{n/m})\sum_{j\in S\cup B^{\prime}}p_{j}$
$\leq O(\sqrt{n/m})\cdot\textup{{OPT}}.$
The last inequality holds because all jobs in $S\cup U^{\prime}$ can be charged to a corresponding larger job in the original job set $J$. Notably, a proxy job appears only when a larger job arrives, and thus, we can charge it to that larger job. Finally, since our algorithm schedules $S\cup U\cup B$ exactly the same as $\textup{{NSJF}}(S\cup U^{\prime}\cup B^{\prime})$, the lemma follows. ∎
@@
Let $B$ be the set of blocked small jobs. Then the total flow time of jobs in $B$ is bounded by:
$\sum_{j\in B}F_{j}\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
@@
Consider an arbitrary blocked small job $j$. By our scheduling rule, $j$ can only experience waiting time when all Small-Only machines are busy, since $j$ is classified as a small job. Therefore, during the entire interval $r_{j},s_{j})$, all Small-Only machines must be fully occupied. We refer to this as a busy time of Small-Only machines.
On the other hand, due to the restart rule, at any busy time $t$ of Small-Only machines, the number of blocked small jobs still waiting is at most $\sqrt{nm}$. This is because the number of blocked small jobs increases only when a new large job starts processing. However, when that happens, all previously blocked small jobs are scheduled and no longer waiting.
@@
We remark that when $m=1$, there is only one Small-Only machine, and this argument no longer holds. Finally, the total flow time of all blocked small jobs is:
$\sum_{j\in B}F_{j}=\sum_{j\in B}(s_{j}-r_{j}+p_{j})\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
For each large job $j$, consider its completion time $C_{j}$, which is either its own completion time or the completion time of its proxy job $j^{\prime}$. If it is proxied, the delay from $r_{j^{\prime}}$ to $C_{j^{\prime}}=C_{j}$ is already charged to the small job flow time in Lemma 12 and Lemma 13. Thus, we only need to analyze the modified flow time where
@@
For large job set $L$. We have $\sum_{j\in L}\hat{F}_{j}\leq O(\sqrt{n/m})\cdot\textup{{OPT}}$.
@@
Ignoring the final $p_{j}$ units of processing time, we focus on the delay interval $r_{j},s_{j})$ (or $r_{j},r_{j^{\prime}})$ if $j$ is proxied by $j^{\prime}$). Define $t_{j}=s_{j}$ (or $t_{j}=r_{j^{\prime}}$ if proxied), and aim to bound $t_{j}-r_{j}$.
We begin by showing that shifting the release time $r_{j}$ by $\sqrt{n/m}\cdot p_{j}$ is acceptable. Let $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}$. Then:
@@
  1. 1.
    Genuine busy time: All Mixed machines are busy at $t$ originally, and at least $\lceil m/4\rceil$ of the running jobs are not restarted later.
  2. 2.
    Fake busy time: All Mixed machines are busy at $t$ originally, but more than $\lceil m/4\rceil$ of the running jobs are restarted.
We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts: $F_{j}^{(1)}$ and $F_{j}^{(2)}$, corresponding to time spent in genuine busy and fake busy intervals, respectively:
$t_{j}-t^{\prime}_{j}=F_{j}^{(1)}+F_{j}^{(2)}.$
Bounding $F_{j}^{(1)}$: At any genuine busy time, at least $\lceil m/4\rceil$ machines are processing jobs that are not restarted. The total duration of such intervals is at most $O(\textup{{OPT}}/m)$. Furthermore, each such interval can contribute to the delay of at most $\sqrt{nm}$ different large jobs (by Lemma 11). Hence:
$\sum_{j\in L}F_{j}^{(1)}\leq\sqrt{nm}\cdot O\left(\frac{\textup{{OPT}}}{m}\right)=O\left(\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}\right).$ (5)
Bounding $F_{j}^{(2)}$: Next, we analyze the contribution of fake busy intervals, which are induced by kill events. Let the $i$-th kill event occur at time $\beta_{i}$. For a time $t$ to contribute to $F_{j}^{(2)}$, the following conditions must hold:
  1. 1.
    $t>\beta_{i-1}$ and $t>t^{\prime}_{j}$,
  2. 2.
    $t$ is not a genuine busy time.
We now prove that such $t$ satisfies:
$\beta_{i}-t\leq p_{j}.$ (6)
Proof of the claim: Since $t$ is a fake busy time, more than $\lceil m/4\rceil$ jobs are restarted at $\beta_{i}$. Suppose, for contradiction, that $\beta_{i}-t>p_{j}$. Then these jobs must have been running continuously since $t$, implying that their processing times are at least $p_{j}$. By our algorithm, it is impossible that job $j$ is already released when we decide to start one of these larger jobs. Therefore, letting $U$ denote this set of jobs, we have:
@@
However, by definition $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}$, and assuming the total workload released up to $r_{j}$ is $P(r_{j})$, we have:
$t^{\prime}_{j}\geq r_{j}+\frac{4P(r_{j})}{m}\geq r_{j}+\frac{4\sum_{u\in U}p_{u}}{m}\geq r_{j}+\min_{u\in U}p_{u}>\beta_{i}.$
This proves (6), by $\beta_{i}-t\leq\beta_{i}-t^{\prime}_{j}<0$.
Thus, for each kill event, the total contribution to $\sum_{j}F_{j}^{(2)}$ is at most:
@@
Finally, we bound the number of kill events. Since each kill is triggered only after at least $\lfloor\sqrt{nm}\rfloor$ new small jobs are released, becoming blocked, the total number of restart events is at most $O(\sqrt{n/m})$. Therefore:
$\sum_{j\in L}F_{j}^{(2)}\leq\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}.$ (7)
@@
$\sum_{j\in L}\hat{F_{j}}\leq\sum_{j\in L}\left(p_{j}+t_{j}-t^{\prime}_{j}+t^{\prime}_{j}-r_{j}\right)\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
@@
There exists an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization, e.g., Algorithm 4, that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution, provided that $m\geq 2$.
@@
$\sum_{j\in J}F_{j}\leq\sum_{j\in S}F_{j}+\sum_{j\in B}F_{j}+\sum_{j\in L}\hat{F}_{j}\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$

### v3

@@ 6 Deterministic Algorithm with Kill-and-Restart
In the absence of blocked small jobs and committed large jobs, the total flow time contributed by the jobs in $S$ is simply $F(\textup{{NSJF}}(S))$, which is at most $O\!\left(\sqrt{n/m}\right)\cdot F(\textup{{OPT}}(S))$ by Lemma 2. Indeed, active large jobs never interfere with the scheduling of unblocked small jobs: if a job $j$ is unblocked, then no active large job is running at its release time, and no active large job can be started before $j$, since every such job is larger than $j$.
We now account for the effect of blocked small jobs and committed large jobs.
To analyze their effect, we partition the time line into intervals
$I_{h}=[t_{h}^{(\ell)},t_{h}^{(r)}),\qquad h=1,2,\dots,H,$
defined recursively as follows. We let $t_{1}^{(\ell)}$ be the first time at which an unblocked small job is released. Given $t_{h}^{(\ell)}$, we let $t_{h}^{(r)}$ be the first time at or after $t_{h}^{(\ell)}$ at which the algorithm starts processing an active large job, which may later be killed or retired. If no such time exists, we set $t_{h}^{(r)}=\infty$ and the construction completes. Otherwise, we let $t_{h+1}^{(\ell)}$ be the first time after $t_{h}^{(r)}$ at which an unblocked small job is released, and continue recursively. By construction, every unblocked small job is released in some interval $I_{h}$.
The key properties of each interval $I_{h}$ are the following. First, no active large job is processed inside $I_{h}$: by definition, no large job is processing at $t_{h}^{(\ell)}$ since an unblocked small job is released at $t_{h}^{(\ell)}$, and no large job starts processing inside $I_{h}$. Second, every blocked small job released before $t_{h}^{(\ell)}$ must start processing before $t_{h}^{(r)}$: the algorithm prioritizes small jobs over the large job starting at $t_{h}^{(r)}$.
For each $h$, let $S_{h}$ be the set of unblocked small jobs released in $I_{h}$, and let $K_{h}$ be the set of blocked small jobs that started processing in $I_{h}$. For every job in $K_{h}$, we modify its release time to its start time and obtain the corresponding set $K_{h}^{\prime}$.
Shifting the time origin to $t_{h}^{(\ell)}$, the schedule of $S_{h}\cup K_{h}^{\prime}$ inside $I_{h}$ can be viewed as $\textup{{NSJF}}(S_{h}\cup K_{h}^{\prime},\vec{b_{h}})$, where $\vec{b_{h}}$ is the initial blocking vector: machine $i$ is unavailable during $0,b_{h,i})$ due to a small job or a committed large job still being processed at time $t_{h}^{(\ell)}$, with $b_{h,i}=0$ if machine $i$ is idle. Therefore,
$\sum_{j\in S_{h}}F_{j}\leq\sum_{j\in S_{h}\cup K^{\prime}_{h}}F_{j}=F(\textup{{NSJF}}(S_{h}\cup K_{h}^{\prime},\vec{b_{h}})).$
We now apply the generalized NSJF bound from Lemma 29. Since every small job or committed large job has size at most $O(\textup{{OPT}}/\sqrt{nm})$, the total blocking time $B_{h}=\sum_{i}b_{h,i}$ satisfies $B_{h}\leq m\cdot O(\textup{{OPT}}/\sqrt{nm})=O(\sqrt{m/n}\cdot\textup{{OPT}})$. Hence,
$F(\textup{{NSJF}}(S_{h}\cup K_{h}^{\prime},\vec{b_{h}}))\leq F(\textup{{OPT}}^{m}(S_{h}\cup K_{h}^{\prime}))+O\!\left(\frac{|S_{h}|+|K_{h}^{\prime}|}{\sqrt{nm}}\right)\textup{{OPT}},$
where $\textup{{OPT}}^{m}$ denotes the optimal preemptive schedule with migration.
@@
It remains to compare $F(\textup{{OPT}}^{m}(S_{h}\cup K_{h}^{\prime}))$ with $F(\textup{{OPT}}(S_{h}))$. For this, consider an optimal preemptive schedule $\textup{{OPT}}(S_{h})$. We insert the jobs in $K_{h}^{\prime}$ into $\textup{{OPT}}(S_{h})$ one by one, in a migratory fashion: for each job $j\in K_{h}^{\prime}$, we scan forward from $r_{j}$ and process $j$ at the first available idle slot without delaying the original schedule. If $j$ overlaps a busy period, we interrupt it and resume it at the next idle slot, possibly on a different machine.
Let $F_{j}^{\prime}$ denote the resulting flow time of job $j\in K_{h}^{\prime}$ in this schedule. Since $j$ is only delayed during busy periods, its waiting time $F_{j}^{\prime}-p_{j}$ is at most the total busy time, which is $\sum_{i\in S_{h}\cup K_{h}^{\prime}}p_{i}/m$.
It remains to bound $|K_{h}^{\prime}|$. By the restart rule, if a job belongs to $K_{h}$, then it must have been released while some machine was processing a currently unproxied large job, and hence before $t_{h}^{(\ell)}$. Therefore, every job in $K_{h}$ is already waiting at time $t_{h}^{(\ell)}$. By Lemma 12, the number of waiting blocked small jobs is at most $\sqrt{nm}$, so
$|K_{h}|=|K_{h}^{\prime}|\leq\sqrt{nm}.$
It follows that
$F(\textup{{OPT}}^{m}(S_{h}\cup K_{h}^{\prime}))\leq F(\textup{{OPT}}(S_{h}))+O\!\left(\sqrt{\frac{n}{m}}\right)\sum_{j\in S_{h}\cup K_{h}^{\prime}}p_{j}.$
Combining the above bounds and summing over all $h$, we get
$\sum_{j\in S}F_{j}$ $\leq\sum_{h=1}^{H}F(\textup{{NSJF}}(S_{h}\cup K_{h}^{\prime},\vec{b_{h}}))$
$\leq\sum_{h=1}^{H}F(\textup{{OPT}}^{m}(S_{h}\cup K_{h}^{\prime}))+O\!\left(\sum_{h=1}^{H}\frac{|S_{h}|+|K_{h}^{\prime}|}{\sqrt{nm}}\right)\textup{{OPT}}$
$\leq\sum_{h=1}^{H}F(\textup{{OPT}}(S_{h}))+O\!\left(\sum_{h=1}^{H}\frac{|S_{h}|+|K_{h}^{\prime}|}{\sqrt{nm}}\right)\textup{{OPT}}$
$\quad+O\!\left(\sqrt{\frac{n}{m}}\right)\sum_{h=1}^{H}\sum_{j\in S_{h}\cup K_{h}^{\prime}}p_{j}.$
Finally, the sets $\{S_{h}\}_{h=1}^{H}$ partition $S$, and the sets $\{K_{h}^{\prime}\}_{h=1}^{H}$ partition $K^{\prime}$. Thus,
$\sum_{h=1}^{H}F(\textup{{OPT}}(S_{h}))\leq F(\textup{{OPT}}(S)),\qquad\sum_{h=1}^{H}(|S_{h}|+|K_{h}^{\prime}|)\leq n.$
The total processing volume of the jobs in $\bigcup_{h}(S_{h}\cup K_{h}^{\prime})$ is at most that of the original jobs in $J$, giving
$\sum_{h=1}^{H}\sum_{j\in S_{h}\cup K_{h}^{\prime}}p_{j}\leq O(\textup{{OPT}}).$
Therefore,
$\sum_{j\in S}F_{j}\leq O\!\left(\sqrt{\frac{n}{m}}\right)\textup{{OPT}}.$
@@
Let $K$ be the set of blocked small jobs. Then the total flow time of jobs in $K$ is bounded by:
$\sum_{j\in K}F_{j}\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
@@
Consider an arbitrary blocked small job $j$. By our scheduling rule, $j$ can only experience waiting time when all Small-Only machines are busy, since $j$ is classified as a small job. During the entire interval $r_{j},s_{j})$, all Small-Only machines must be fully occupied. We refer to this as a busy time of Small-Only machines.
Due to the restart rule, at any busy time $t$ of Small-Only machines, the number of blocked small jobs still waiting is at most $\sqrt{nm}$ by Lemma 12.
@@
We remark that when $m=1$, there is no Small-Only machine, and this argument no longer holds. Finally, the total flow time of all blocked small jobs is:
$\sum_{j\in K}F_{j}=\sum_{j\in K}(s_{j}-r_{j}+p_{j})\leq O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
For each large job $j$, we only need to analyze the modified flow time:
@@
For the large-job set $L$, we have $\sum_{j\in L}\hat{F}_{j}\leq O(\sqrt{n/m})\cdot\textup{{OPT}}$.
@@
We focus on the delay interval $r_{j},t_{j})$, where
$t_{j}=\begin{cases}r_{j^{\prime}}&\text{if $j$ is proxied by $j^{\prime}$,}\\ s_{j}&\text{otherwise,}\end{cases}$
and aim to bound $t_{j}-r_{j}$. We begin by showing that shifting the release time $r_{j}$ by $\sqrt{n/m}\cdot p_{j}$ is acceptable. Let $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}$. Then:
@@
  1. $(\checkmark)$
    Genuine busy time: All Mixed machines are busy at $t$, and at least $\lceil m/4\rceil$ of the jobs running on Mixed machines are not restarted later.
  2. $(\times)$
    Fake busy time: All Mixed machines are busy at $t$, but more than $\lceil m/4\rceil$ of the jobs running on Mixed machines will be restarted later.
We decompose the delay $t_{j}-t^{\prime}_{j}$ into two parts: $F_{j}^{(\checkmark)}$ and $F_{j}^{(\times)}$, corresponding to time spent in genuine busy and fake busy intervals, respectively:
$t_{j}-t^{\prime}_{j}=F_{j}^{(\checkmark)}+F_{j}^{(\times)}.$
Bounding $F_{j}^{(\checkmark)}$ (genuine busy).
At any genuine busy time, at least $\lceil m/4\rceil$ machines are processing jobs that are not restarted. The total duration of such intervals is at most $O(\textup{{OPT}}/m)$. Since there are at most $\ell=\lfloor\sqrt{nm}\rfloor$ large jobs at any time by Lemma 11, we have:
$\sum_{j\in L}F_{j}^{(\checkmark)}\leq\sqrt{nm}\cdot O\left(\frac{\textup{{OPT}}}{m}\right)=O\left(\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}\right).$ (5)
Bounding $F_{j}^{(\times)}$ (fake busy).
Next, we analyze the contribution of fake busy intervals induced by the kill events. Let the $i$-th kill event occur at time $\beta_{i}$. The key observation is the following.
Claim 1.
For any $t\in[t^{\prime}_{j},t_{j})$ that is a fake busy time between consecutive kill events $\beta_{i-1}$ and $\beta_{i}$, we have $\beta_{i}-t\leq p_{j}$.
Proof of 1.
Since $t$ is a fake busy time, more than $\lceil m/4\rceil$ jobs are restarted at $\beta_{i}$. Suppose, for contradiction, that $\beta_{i}-t>p_{j}$. Then these jobs must have been running continuously since $t$, implying that their processing times are at least $p_{j}$. By our algorithm, job $j$ had not yet been released when the algorithm started one of these larger jobs. Therefore, letting $U$ denote this set of jobs, we have:
@@ 2. $(\times)$
However, by the partition rule $p_{j}\geq 4P(r_{j})/\ell$ and $\ell=\lfloor\sqrt{nm}\rfloor$, we have $t^{\prime}_{j}=r_{j}+\sqrt{n/m}\cdot p_{j}\geq r_{j}+4P(r_{j})/m$. Since the jobs in $U$ were released before $r_{j}$, their total size is at most $P(r_{j})$. Moreover, $|U|>\lceil m/4\rceil$, so:
$\frac{4\sum_{u\in U}p_{u}}{m}\geq\frac{4\cdot|U|\cdot\min_{u\in U}p_{u}}{m}>\min_{u\in U}p_{u}.$
Combining,
$t^{\prime}_{j}\geq r_{j}+4P(r_{j})/m\geq r_{j}+4\sum_{u\in U}p_{u}/m>r_{j}+\min_{u\in U}p_{u}>\beta_{i}.$
This gives $\beta_{i}-t\leq\beta_{i}-t^{\prime}_{j}<0$, a contradiction. ∎
By 1, each large job $j$ contributes at most $p_{j}$ to $F_{j}^{(\times)}$ per kill event. Thus, for each kill event, the total contribution to $\sum_{j}F_{j}^{(\times)}$ is at most:
@@
Finally, we bound the number of kill events. Since each kill is triggered only after at least $\lfloor\sqrt{nm}\rfloor$ new blocked small jobs are released, the total number of kill events is at most $O(\sqrt{n/m})$. Therefore:
$\sum_{j\in L}F_{j}^{(\times)}\leq\sqrt{\frac{n}{m}}\cdot\textup{{OPT}}.$ (6)
@@
$\sum_{j\in L}\hat{F}_{j}\leq\sum_{j\in L}\left(p_{j}+t_{j}-t^{\prime}_{j}+t^{\prime}_{j}-r_{j}\right)=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$
@@
Algorithm 4 is an online polynomial-time deterministic algorithm with kill-and-restart for total flow time minimization that is $O(\sqrt{n/m})$-competitive against the preemptive offline solution, provided that $m\geq 2$.
@@
$\sum_{j\in J}F_{j}\leq\sum_{j\in S}F_{j}+\sum_{j\in K}F_{j}+\sum_{j\in L}\hat{F}_{j}=O\left(\sqrt{\frac{n}{m}}\right)\cdot\textup{{OPT}}.$